### v1

# GS4: Generalizable Sparse Splatting Semantic SLAM

Abstract
Traditional SLAM algorithms are excellent at camera tracking but might generate lower resolution and incomplete 3D maps. Recently, Gaussian Splatting (GS) approaches have emerged as an option for SLAM with accurate, dense 3D map building. However, existing GS-based SLAM methods rely on per-scene optimization which is time-consuming and does not generalize to diverse scenes well. In this work, we introduce the first generalizable GS-based semantic SLAM algorithm that incrementally builds and updates a 3D scene representation from an RGB-D video stream using a learned generalizable network. Our approach starts from an RGB-D image recognition backbone to predict the Gaussian parameters from every downsampled and backprojected image location. Additionally, we seamlessly integrate 3D semantic segmentation into our GS framework, bridging 3D mapping and recognition through a shared backbone. To correct localization drifting and floaters, we propose to optimize the GS for only 1 iteration following global localization. We demonstrate state-of-the-art semantic SLAM performance on the real-world benchmark ScanNet with an order of magnitude fewer Gaussians compared to other recent GS-based methods, and showcase our model’s generalization capability through zero-shot transfer to the NYUv2 and TUM RGB-D datasets.
1 Introduction
Semantic Visual Simultaneous Localization and Mapping (SLAM) is a long-standing challenge in computer vision that aims to reconstruct a dense 3D semantic map of an environment while simultaneously estimating camera poses from an input video. These dense semantic maps have widespread applications in autonomous driving, AR/VR and robot navigation, enabling robots to understand and interact with their surroundings more effectively by providing rich 3D spatial and semantic information.
Traditional visual SLAM systems consist of several independent components, including keypoint detection, matching, and bundle adjustment [25]. Although traditional SLAM systems generally provide accurate localization, they struggle to generate accurate, dense 3D maps, limiting their effectiveness in robotics applications that require precise environmental interaction, such as manipulation tasks. Recent advances in differentiable rendering [23, 14] have opened up new opportunities for more streamlined yet more powerful visual SLAM approaches. Neural scene representations, such as Neural Radiance Fields (NeRF) [23], have been shown effective in generating accurate, dense 3D maps and have been successfully adopted as a 3D scene representation for visual SLAM [35, 45, 11]. However, volumetric rendering used in NeRF remains computationally expensive because a neural network needs to be trained on-the-fly, which leads to a compromise between reconstruction quality and training costs.
Recently, 3D Gaussians have emerged as a promising 3D scene representation, offering fast, differentiable, and high quality rendering capabilities [15]. With these advantages, Gaussian-based representations have been shown to be well-suited for SLAM systems [13, 22]. However, existing approaches require test-time gradient-based optimization to estimate 3D Gaussians for each scene independently, which is computationally expensive and unsuitable for real-time applications. Moreover, these methods rely on heuristic-based Gaussian densification and pruning [15], often leading to dense Gaussian representations that do not scale well in large, real-world environments.
In this paper, we propose a novel GS-based SLAM system GS4 (Generalizable Sparse Splatting Semantic SLAM), based on the idea of directly predicting 3D semantic Gaussians with a learned feed-forward network. With this approach, we no longer require the expensive test-time network learning or per-scene optimization. Besides, by incorporating an image recognition backbone, we can utilize the same features to predict both the Gaussian parameters as well as the semantics of each Gaussian, creating a unified approach for semantic SLAM without incorporating an external semantic segmentation algorithm.
GS4 directly predicts 3D semantic Gaussians from RGB-D images using a learned feed-forward network. Additionally, GS4 features a Gaussian refinement network that effectively merges Gaussians predicted from different video frames, replacing the handcrafted heuristics used for Gaussian densification and pruning in prior GS-based methods. Our refinement network enables a learned pruning strategy, producing a compact 3D Gaussian representation with an order of magnitude fewer Gaussians than competiting approaches. To adapt to global localization (bundle adjustment) schemes that would change past camera poses and point locations, we propose to optimize Gaussian parameters for 1 iteration after global localization. This allowed the algorithm to correct the Gaussian parameters while alleviating the common issue of “floaters” in feed-forward GS methods that float in the air and reduce map quality.
We demonstrate that GS4 achieves state-of-the-art performance across all key metrics in localization, mapping, and segmentation on the real-world benchmark ScanNet, while using only of Gaussians compared to prior GS SLAM methods (Fig. 1). Furthermore, we highlight the generalization capability of our system via zero-shot transfer to the NYUv2 and TUM RGB-D datasets, which, to the best of our knowledge, is the first demonstration of semantic SLAM generalization in a modern neural SLAM system.
In summary, our contributions are as follows:
We propose GS4, the first generalizable Gaussian splatting semantic SLAM approach on monocular RGB-D sequences. Results showed that GS4 obtains state-of-the-art on real ScanNet scenes and also zero-shot generalizes to the real NYUv2 and TUM RGB-D datasets without any fine-tuning.
Our proposed Gaussian refinement network effectively merges Gaussians from different frames into a 3D representation, while significantly reducing the number of Gaussians required to represent a scene.
Our proposed 1-iteration Gaussian optimization significantly improves reconstruction quality with a small additional computational cost.
2 Related Work
Traditional SLAM: Early visual SLAM methods [25] demonstrated robust localization through effective keypoint detection and matching, which resulted in sparse 3D reconstructions. While these approaches provided reliable localization, the sparse nature of the reconstructed maps limited their utility in applications requiring detailed 3D maps. To address this issue, dense visual SLAM [16, 6] focused on constructing detailed maps to support applications like augmented reality (AR) and robotics. Prior methods [2, 8, 27, 1, 38, 29] achieved dense mapping by employing representations based on Signed Distance Fields (SDF), rather than relying on sparse representations such as point clouds or grids. However, these approaches often suffer from over-smoothed reconstruction, failing to capture fine details crucial for certain tasks.
NeRF-based SLAM: Neural Radiance Fields (NeRF) [23] has gained popularity as a 3D scene representation due to its ability to generate accurate and dense reconstructions. NeRF employs Multi-Layer Perceptron (MLP) to encode scene information and performs volume rendering by querying opacity and color along pixel rays. Methods such as iMAP [35], NICE-SLAM [45], and ESLAM [11] incorporate this implicit scene representation into SLAM, leveraging NeRF’s high-fidelity reconstructions to improve both localization and mapping. However, the volumentric rendering process in NeRF is computationally expensive, often requiring trade-offs such as limiting the number of pixels during rendering, These trade-offs, while improving efficiency, may compromise the system’s accuracy in both localization and mapping.
GS-based SLAM: 3D Gaussian Splatting (3DGS) [15] employs splatting rasterization instead of ray marching. This approach iterates over 3D Gaussian primitives rather than marching along rays, resulting in a more expressive and efficient representation capable of capturing high-fidelity 3D scenes with significantly faster rendering times. With the efficiency and rapid rasterization capabilities of 3D Gaussian Splatting, GS-based SLAM systems achieve improved accuracy and speed in dense scene reconstruction. SplaTAM [12] introduces silhouette-guided rendering to facilitate structured map expansion, enabling efficient dense visual SLAM. Gaussian Splatting SLAM [22] integrates novel Gaussian insertion and pruning techniques. Additionally, GS-ICP SLAM [9] combines ICP with 3DGS, achieving both faster computational speed and superior map quality. Expanding upon these advancements, SGS-SLAM [18] extends 3D Gaussian representations to include semantic scene understanding. However, existing GS-based SLAM methods employ per-scene optimization, requiring iterative refinement of Gaussians initialized from keyframes through rendering supervision. As a result, SGS-SLAM relies on additional segmentation models to predict semantic labels for each image, creating computational overhead.
Feed-forward Models for GS: Recent research has introduced feed-forward approaches for scene-level 3DGS reconstruction using generalizable models [3, 4, 19]. Unlike previous methods, GS-LRM [40] avoids specialized 3D structural designs, instead uses an attention-based transformer to achieve state-of-the-art results. However, to the best of our knowledge, feed-forward models have been applied on a small number of images and have not yet been introduced in GS-based semantic SLAM approaches.
3 Methods
In this section, we describe our proposed SLAM approach. We first provide a brief overview of Gaussian Splatting, then detail our Gaussian prediction network and Gaussian refinement network. Finally, we explain how these networks are utilized within the entire SLAM system.
3.1 Gaussian Splatting
We represent a 3D map using a set of anisotropic 3D Gaussians used in Gaussian Splatting. Each Gaussian is characterized by RGB color , center position , scale , quaternion , opacity and semantic class vector , where N is the number of classes.
The rendering process is defined as:
where is a quantity of a pixel to be rendered, which can be color, depth or semantic label, and is that quantity of the -th 3D Gaussian, while
defines the visibility of the -th splatted Gaussian. Here, denotes the offset between the pixel coordinates and the 2D mean of the -th Gaussian. Following [13], We also render a silhouette image to determine visibility:
3.2 Gaussian Prediction and Refinement
Our proposed Gaussian prediction network takes RGB-D images as input and predicts 3D Gaussian parameters. These parameters also include semantic labels, enabling the rendering of photometric, geometric, and semantic views. Next, the Gaussian refinement network processes Gaussians predicted from different video frames and learns to merge them, constructing a consistent 3D scene representation.
3.2.1 Backbone for Gaussian Prediction
We train a transformer model to regress 3D GS parameters from an image with a known camera pose (from tracking, described in Sec. 3.3.1), while simultaneously assigning semantic labels to these 3D Gaussians. We start with ODIN [10], a model that starts from multi-view RGB-D images with a transformer architecture that alternates between 2D within-view and 3D cross-view attention for information fusion over multiple views. To further enhance efficiency, we have made the following modifications to the ODIN architecture:
We modify the ODIN architecture by removing the 3D attention layers between all stages of the 2D backbone, retaining only the 3D attention layers after their multi-scale deformable attention layer. This adjustment reduces computational costs while maintaining performance.
ODIN originally utilizes voxel-based operations, such as mean pooling of features and XYZ coordinates within occupied voxels in 3D space, which are beneficial for dense point clouds but can be computationally intensive. Given that our inference involves single-view input (i.e., sparser point cloud) at a time, we eliminate these voxel-based operations, resulting in faster processing without compromising performance.
ODIN utilizes 3D interpolation to upsample low-resolution feature maps. We replace this step with the point-based upsampling layer from AutoFocusFormer [46], which is more efficient.
These modifications collectively enhance the efficiency of the ODIN architecture ( speedup compared to the original), making it more suitable for our SLAM system.
Generating Prediction Tokens with ODIN. We concatenate an RGB image and a depth image resulting in a 4-channel feature map that is fed into the model:
where denotes the total number of prediction tokens per image. The variable represents the network level at which Gaussians are predicted. The backbone consists four levels, each processing a progressively downsampled set of tokens. If the second level is chosen, the feature map has a spatial resolution of , resulting in tokens per image. Here, represents the encoder features of an -th token from the backbone, while denotes the corresponding decoder features at the same resolution.
Processing Prediction Tokens with Transformer. Given the selected prediction level, we concatenate the encoder features and decoder features for each token , and process them through local-attention transformer layers [46]: to obtain the final features for the -th token, integrating information from both the encoder and decoder.
Decoding Prediction Tokens to Gaussians. Each output token’s features, , from the transformer layers are decoded into Gaussian parameters using Multi-Layer Perceptron (MLP):
Here, and represent the offsets for the 2D position (, ) of Gaussian in image space, while is the offset for the noisy depth obtained from the depth image. These offsets are added to the original values, which are then back-projected into 3D space using the camera’s intrinsic and extrinsic parameters, yielding the 3D center position . Similarly, represents the offset for the RGB value, which is added to the original RGB value observed from the image, yielding the final RGB value for each Gaussian.
Besides Gaussian parameters, the mask decoder head in ODIN predicts segmentation labels for the input image. By indexing these labels based on the tokens’ positions, we assign a semantic class vector to each Gaussian. Finally, each Gaussian is also assigned the final feature vector of its corresponding token for use in the subsequent Gaussian refinement stage.
To supervise the prediction of semantic segmentation, we follow the setup in Mask2Former [5]. We denote the corresponding segmentation loss by . In addition, we render images at supervision views—comprising the current input view and randomly selected novel views that overlap with the current input—using the predicted Gaussians from the current input, and minimize RGB-D and semantic rendering loss. For novel view supervision, we focus solely on areas visible in the input view, ensuring that the optimization process focuses on regions consistently observed across both input and novel views. We explain the loss functions used during training below.
RGB Rendering Loss. Following previous work [40, 47], we use a combination of the Mean Squared Error (MSE) loss and Perceptual loss:
where is the weight for the perceptual loss.
Depth Rendering Loss. For depth images, we use L1 loss:
Semantic Rendering Loss. For semantic rendering, we use the cross entropy loss:
where the rendered semantic image has N channels, each corresponding to a different semantic category.
Overall Training Loss. Our total loss comprises multiple rendering losses and the segmentation loss :
where we use , and .
3.2.2 Gaussian Refinement Network
In our SLAM system, we merge predicted Gaussians from a new frame into the 3D map. To ensure that the combined set of Gaussians accurately represents the scene, we propose a novel Gaussian Refinement Network to refine both the existing Gaussians in the 3D map and the newly added ones, enabling their effective merging. The input to the network includes the features and 3D positions of all Gaussians from both the 3D map and the new frame. We process these using several local-attention transformer layers with 3D neighborhoods to fuse and update the features for each Gaussian. Subsequently, MLP layers predict updates , , and for each Gaussian. These updates refine the Gaussians to accurately render both current and previous views. To supervise the network, we render the current input view along with overlapping previous views. The total training loss is defined as follows:
where we use , and .
During testing time, we introduce a threshold to manage the uncertainty of each Gaussian. Once a Gaussian has been updated times by the refinement network, we consider its uncertainty sufficiently reduced and exclude them from further updates. We set in our experiments. After Gaussian refinement, we prune Gaussians whose updated opacity falls below 0.005, effectively removing those that have become unimportant after merging.
3.3 The SLAM System
An overview of the system is summarized in Fig. 2. The Gaussian prediction network predicts 3D Gaussians from RGB-D images, which can be rendered into high-fidelity color, depth, and semantic images. The Gaussian refinement network refines both the existing Gaussians in the 3D map and the newly added ones to accurately render both current and previous views. During testing, we occasionally run 1-iteration test-time optimization and refine 3D Gaussians in the map to reflect camera pose updates from loop closure and bundle adjustment in the tracking module.
3.3.1 Tracking and Global Bundle Adjustment
For camera tacking in our SLAM system, we adopt a tracker used in GO-SLAM [42] which is an enhanced version of DROID-SLAM’s tracking module [36]. In the Local Camera Tracking, a keyframe is initialized when sufficient motion is detected, and loop closure (LC) is performed. Meanwhile, the Global Localization is equipped with full bundle adjustment (BA) for real-time global refinement. Both LC and BA help address the problem of accumulated errors and drift that can occur during the localization process.
Local Camera Tracking. Local camera tracking processes RGB-D images to compute optical flow between each new frame and the previous keyframe [36]. If the average flow exceeds a predefined threshold, the current frame is selected as a new keyframe and added to the keyframe buffer for subsequent refinement. In the keyframe buffer, a loop candidate is accepted only if a frame pair mean flow is below a predefined threshold and the frames are far apart enough. Subsequently, a differentiable Dense Bundle Adjustment (DBA) [36] layer is applied to refine the camera pose for each keyframe.
Global Localization. Full bundle adjustment (BA) is performed in a separate thread, following strategies employed in [24]. This approach allows the system to continue tracking new frames and handling loop closures concurrently.
3.3.2 One-Iteration Gaussian Optimization
Loop closure and bundle adjustment are essential components in SLAM systems, employed to correct accumulated drift and adjust the camera poses of previous frames. However, these adjustments can cause Gaussians inserted based on earlier, uncorrected poses to misalign with the scene, leading to inaccurate rendering and mapping. It is crucial to implement a mechanism that updates the Gaussians in the 3D map following pose corrections. To address this issue, we propose using rendering-based optimization to update the Gaussian parameters , , and with only a single iteration. We render RGB-D images for the top-k frames, selected based on significant pose changes. This approach maintains the consistency of the 3D map with updated camera poses. To enhance the efficiency of this optimization, we employ the batch rendering technique from [39]. We omit semantic image rendering to improve system efficiency. For 1-iteration optimization, we add a SSIM term to the RGB loss, following [15]:
where is set to for all experiments.
4 Experiments
4.1 Experimental Setup
Training Settings. We train our Gaussian prediction and refinement networks entirely on RGB-D videos from the real ScanNet dataset. Excluding the six test scenarios commonly used in SLAM benchmarks, we utilize all remaining train and val scenarios for training, employing 20 common semantic classes for supervision. We use AutoFocusFormer [46] as the backbone for ODIN and for Gaussian prediction, we select the second stage of the backbone as the prediction stage. To train the Gaussian prediction network, we adopt a low-to-high resolution curriculum as described in [47], progressing through three stages with image resolutions of 256×256, 480×480, and 640×480. In the first two stages, training images are resized so that the shorter side is 256 and 480 pixels, respectively, and then center-cropped to a square. For the Gaussian refinement network, which operates across multiple consecutive frames, we use a multi-stage training strategy: starting with two consecutive frames, progressing to four, and finally using eight.
Datasets. During testing, we evaluate our method on six real-world scenes on ScanNet [7]. Additionally, we perform zero-shot experiments on several real scenes from NYUv2 [26] and TUM RGB-D [34], as well as synthetic scenes from the Replica dataset [32].
Metrics. We use PSNR, Depth-L1 [44], SSIM [37], and LPIPS [41] to evaluate the reconstruction and rendering quality. We additionally report reconstruction metrics such as Accuracy, Completion, Completion Ratio (7cm) and F1 (7cm) in the supplemetary. For GS-based SLAM methods, we also show the number of Gaussians for each scene. For semantic segmentation, we report the mean Intersection over Union (mIoU). To evaluate the accuracy of the camera pose, we adopt the average absolute trajectory error (ATE RMSE) [33].
Baselines. We compare our method to several state-of-the-art approaches, including NeRF-based SLAM methods such as NICE-SLAM [44], GO-SLAM [42] and Point-SLAM [31], as well as 3D Gaussian-based SLAM methods like SplaTAM [13], RTG-SLAM [28], GS-ICP SLAM [9], and Semantic 3D Gaussian-based SLAM such as SGS SLAM [18] and OVO-Gaussian-SLAM [21]. SGS-SLAM and OVO-Gaussian-SLAM are the only semantic SLAM methods available for comparison since the code is not available for other semantic SLAM approaches. Note that SGS-SLAM [18] and DNS-SLAM [17] employ test-time optimization using ground truth semantic labels on the testing set. And SGS-SLAM has been shown to outperform all other existing semantic SLAM methods [43, 17]. To ensure a fair comparison and simulating SLAM applications in real-world scenarios where ground truth semantic labels are unavailable, we trained a 2D segmentation model using a Swin backbone [20] with Mask2Former [5] on ScanNet, following the same training strategy as our model, and used predicted semantic labels to supervise SGS-SLAM.
4.2 Results
| Methods | Metrics | 0000 | 0059 | 0106 | 0169 | 0181 | 0207 | Avg |
| PSNR | 18.71 | 16.55 | 17.29 | 18.75 | 15.56 | 18.38 | 17.54 | |
| NICE-SLAM | SSIM | 0.641 | 0.605 | 0.646 | 0.629 | 0.562 | 0.646 | 0.621 |
| LPIPS | 0.561 | 0.534 | 0.510 | 0.534 | 0.602 | 0.552 | 0.548 | |
[…]
| Point-SLAM | SSIM | 0.806 | 0.765 | 0.676 | 0.686 | 0.823 | 0.750 | 0.751 |
| LPIPS | 0.485 | 0.499 | 0.544 | 0.542 | 0.471 | 0.544 | 0.514 | |
[…]
| SSIM | 0.589 | 0.792 | 0.748 | 0.776 | 0.615 | 0.694 | 0.702 | |
| SplaTAM | LPIPS | 0.485 | 0.289 | 0.314 | 0.281 | 0.456 | 0.358 | 0.364 |
| Depth L1 | 8.01 | 5.33 | 7.72 | 6.18 | 8.33 | 7.67 | 7.21 | |
| GS Num | 3089k | 2117k | 1821k | 1656k | 2714k | 3402k | 2466k | |
[…]
| SSIM | 0.375 | 0.291 | 0.401 | 0.408 | 0.390 | 0.367 | 0.372 | |
| RTG SLAM | LPIPS ↓ | 0.788 | 0.780 | 0.716 | 0.711 | 0.786 | 0.787 | 0.761 |
| Depth L1 | 90.5 | 116.00 | 77.40 | 92.50 | 91.90 | 117.03 | 97.56 | |
| GS Num | 1940k | 1201k | 1000k | 1101k | 1113k | 1021k | 1229k | |
[…]
| SSIM | 0.594 | 0.622 | 0.655 | 0.642 | 0.722 | 0.633 | 0.645 | |
| GS-ICP SLAM | LPIPS ↓ | 0.728 | 0.703 | 0.655 | 0.698 | 0.696 | 0.624 | 0.684 |
| Depth L1 | 218k | 259k | 236k | 223k | 216k | 196k | 224k | |
| GS Num | 5666k | 2586k | 2506k | 2190k | 2519k | 3027k | 2565k | |
[…]
| SSIM | 0.511 | 0.742 | 0.636 | 0.572 | 0.531 | 0.569 | 0.594 | |
| SGS SLAM | LPIPS | 0.687 | 0.485 | 0.597 | 0.613 | 0.706 | 0.600 | 0.615 |
| Depth L1 | 9.11 | 6.73 | 10.66 | 18.89 | 15.02 | 10.56 | 11.83 | |
| GS Num | 2840k | 2020k | 1582k | 1266k | 2054k | 2926k | 2114k | |
[…]
| SSIM | 0.864 | 0.858 | 0.794 | 0.871 | 0.859 | 0.853 | 0.850 | |
| GS4 (Ours) | LPIPS | 0.348 | 0.311 | 0.366 | 0.284 | 0.377 | 0.375 | 0.344 |
| Depth L1 | 3.89 | 6.06 | 9.85 | 7.66 | 4.30 | 4.70 | 6.08 | |
| GS Num | 438k | 227k | 220k | 241k | 246k | 397k | 295k |
| Methods | Metrics | 0000 | 0059 | 0106 | 0169 | 0181 | 0207 | Avg |
| PSNR | 20.76 | 16.24 | 18.11 | 14.93 | 17.99 | 21.20 | 18.21 | |
| SSIM | 0.784 | 0.606 | 0.685 | 0.407 | 0.694 | 0.766 | 0.657 | |
| GO-SLAM | LPIPS | 0.497 | 0.562 | 0.556 | 0.600 | 0.601 | 0.502 | 0.553 |
| Depth L1 | 6.58 | 21.38 | 15.87 | 48.27 | 8.89 | 7.85 | 18.14 | |
[…]
| SSIM | 0.901 | 0.875 | 0.808 | 0.884 | 0.881 | 0.884 | 0.872 | |
| GS4 (Ours) | LPIPS | 0.238 | 0.257 | 0.307 | 0.257 | 0.284 | 0.268 | 0.269 |
| Depth L1 | 3.93 | 6.16 | 9.68 | 7.09 | 4.40 | 4.97 | 6.04 | |
| GS Num | 194k | 99k | 110k | 112k | 127k | 177k | 136k |
GS-ICP SLAM |
SplaTAM |
SGS SLAM |
GS4 (Ours) |
Ground Truth |
| scene0000 | scene0169 | scene0207 | ||||
Rendering and Reconstruction Performance. In Table 4.2, we evaluate the rendering and reconstruction performance of our SLAM method on ScanNet. This is a difficult task compared to the synthetic data where neural RGB-D SLAM methods usually show strong results, because inevitably inaccurate ground truth camera poses and depths make optimization much harder than completely clean synthetic datasets. Compared to existing dense neural RGB-D SLAM methods, our method achieves the best performance in PSNR, SSIM, and Depth L1 metrics. Specifically, our method surpasses the runner-up, Point-SLAM [30], by 2.89 dB in PSNR (a 14.6 percent improvement) and 0.091 in SSIM (a 13.2 percent improvement). Furthermore, our approach utilizes approximately 10x fewer Gaussians than the baselines. This efficiency highlights the effectiveness of our method in achieving high-quality scene representation with reduced computational complexity.
In Table 4.2, we ran our method with input to compare against GO-SLAM which shares the same tracking method as ours but renders at the same low resolution. GS4 maintains the same PSNR and depth prediction quality as its high resolution version and significantly outperforms GO-SLAM across all metrics.
Fig. 3 shows visual results of RGB and depth rendering. Our method demonstrates superior performance than other GS-SLAM methods. Notably, sometimes the depth maps of our approach even outperforms the noisy ground truth depth inputs. For instance, in the first two columns in Fig. 3, our method delivers a more contiguous and complete rendering of the bicycle tires. Similarly, in the middle two columns in Fig. 3, we reconstruct the chair’s backrest nearly entirely, whereas the GT depth data lacks this detail.
| Methods | SGS SLAM | OVO-Gaussian-SLAM | GS4 (Ours) |
| mIoU(2D) | 37.20 | ✗ | 60.40 |
| mIoU(3D) | 18.87 | 32.58 | 50.00 |
Semantic Performance. In Table 3, we present both 2D rendering and 3D mean Intersection over Union (mIoU) scores across the six ScanNet test scenes. For 3D mIoU evaluation, we first align the reconstructed map with the ground-truth mesh and then use 3D neighborhood voting to assign predicted labels. Our method outperforms the previous runner-up, OVO-Gaussian-SLAM, by 17.42 in 3D mIoU. In terms of 2D rendering mIoU, our method surpasses the state-of-the-art semantic GS-based SLAM approach, SGS-SLAM, by 23.2. As illustrated in Fig. 5, our approach achieves superior semantic segmentation accuracy compared to the SGS-SLAM baseline. For example , in the first column of Fig.5, our semantic rendering provides a more accurate representation of the desks, chairs, and night tables than SGS-SLAM.
SplaTAM |
SGS SLAM |
GS4 (Ours) |
GT |
| sudent lounge | dining room | bedroom | ||||
SGS SLAM |
GS4 (Ours) |
GT (labels) |
GT (RGB) |
| scene0000 | scene0169 | scene0207 |
Tracking Performance. Table 4 shows the tracking results for the six test scenes of the ScanNet dataset. Our method uses the same tracking algorithm as GO-SLAM, which is significantly better than other GS-based SLAM methods.
| Methods | 0000 | 0059 | 0106 | 0169 | 0181 | 0207 | Avg |
| NICE-SLAM | 12.00 | 14.00 | 7.90 | 10.90 | 13.40 | 6.20 | 10.70 |
| Point-SLAM | 10.24 | 7.81 | 8.65 | 22.16 | 14.77 | 9.54 | 12.19 |
| SplaTAM | 12.83 | 10.10 | 17.72 | 12.08 | 11.10 | 7.46 | 11.88 |
| RTG SLAM | 191.00 | 191.87 | 126.8 | 97.92 | 112.69 | 146.85 | 144.52 |
| GS-ICP SLAM | 174.68 | NaN | 194.31 | NaN | NaN | 77.64 | – |
| SGS SLAM | 27.53 | 15.51 | 15.38 | 108.38 | 65.38 | 13.82 | 40.97 |
| GO-SLAM | 5.47 | 7.40 | 7.19 | 7.80 | 7.01 | 7.13 | 7.00 |
| GS4 (Ours) | 5.45 | 7.36 | 7.22 | 7.88 | 6.96 | 6.90 | 6.98 |
Runtime Comparison. Table 5 presents a runtime comparison of our method against the baselines at the resolution, conducted on an Nvidia RTX TITAN. FPS of the system is calculated by dividing the total number of frames by the total time to represent the overall system performance. While GS-ICP SLAM is faster than ours, its tracking performance is significantly lower. Our approach is 12x faster than the GS-based SLAM method SplaTAM and 17x faster than the GS-based semantic SLAM method SGS SLAM.
| Methods | Point-SLAM | SplaTAM | RTG-SLAM | GS-ICP SLAM | SGS-SLAM | GS4 (ours) |
| FPS | 0.05 | 0.23 | 1.01 | 3.62 | 0.17 | 2.87 |
Zero-shot Visualization. Fig. 4 illustrates our zero-shot visualization results on the NYUv2 dataset. Despite our models being exclusively trained on the ScanNet dataset, our method demonstrates superior performance on the NYUv2 dataset compared to other GS-based SLAM approaches.
Ablation Study. We conduct an ablation study using all ScanNet test scenes, as shown in Table 6. The results demonstrate that both the Gaussian Refinement Network and the One-Iteration Gaussian Optimization are critical to the performance of GS4. Additionally, Gaussian pruning significantly reduces the number of Gaussians without sacrificing accuracy.
| Design Choice | PSNR [dB] | SSIM | LPIPS | Depth L1 | mIoU | Gs Num |
| GS Prediction | 15.37 | 0.496 | 0.641 | 27.93 | 36.37 | 101k |
| + GS Refinement | 16.78 | 0.586 | 0.565 | 21.66 | 41.58 | 388k |
| + One-Iter. Optimization | 22.69 | 0.849 | 0.346 | 6.10 | 60.01 | 602k |
| + GS Pruning (Full SLAM) | 22.71 | 0.850 | 0.344 | 6.08 | 60.40 | 295 k |
5 Conclusion
We present GS4, a novel SLAM system that incrementally constructs and updates a 3D semantic scene representation from a monocular RGB-D video with a learned generalizable network. Our novel Gaussian refinement network and 1-iteration Gaussian optimization significantly improve the performance of our approach. Our experiments demonstrate state-of-the-art semantic SLAM performance on the ScanNet benchmark while using 10x less Gaussians than baselines. The model also showed strong generalization capabilities through zero-shot transfer to the NYUv2 and TUM RGB-D datasets. In future work, we will further improve the computational speed of GS4 and explore options for a pure RGB-based SLAM approach.
6 Acknowledgements
This work is partially supported by the NSF under grants 1751402 and 2321851, the ONR under grant N0014-21-1-2052, and DARPA under grant HR00112490423. We thank Dr. Stefan Lee for his valuable suggestions and Lianghui Wang and Guowei Li for their assistance with HPC resources.
Appendix
Appendix A Reconstruction Results on ScanNet
In Table 7, we present the reconstruction metrics on the ScanNet dataset. GS4 outperforms all baselines in terms of completion and F-score.
| Methods | Acc. | Comp. | Comp. Ratio (<7cm) | F-Score (<7cm) | GS Num |
| SplaTAM | 8.10 | 5.58 | 76.34 | 75.95 | 2466k |
| RTG-SLAM | 99.80 | 47.44 | 24.61 | 16.69 | 1229k |
| SGS-SLAM | 17.11 | 13.75 | 55.01 | 55.26 | 2114k |
| GS4 | 8.56 | 3.88 | 87.34 | 79.90 | 295k |
Appendix B Zero-Shot Results on NYUv2
In Table B, we present the quantitative zero-shot results across three scenes from the NYUv2 dataset. Our method outperforms all other GS-based SLAM approaches on all rendering metrics, while using significantly fewer Gaussians with much faster processing speed.
| Methods | Metrics | bedroom | student lounge | dining room | Avg |
| PSNR | 17.99 | 20.77 | 17.82 | 18.86 | |
| SSIM | 0.692 | 0.795 | 0.589 | 0.692 | |
| SplaTAM | LPIPS | 0.343 | 0.309 | 0.465 | 0.372 |
| GS Num | 1529k | 1116k | 1063k | 1236k | |
| FPS | 0.32 | 0.33 | 0.33 | 0.33 | |
[…]
| SSIM | 0.146 | 0.299 | 0.217 | 0.221 | |
| RTG-SLAM | LPIPS | 0.738 | 0.662 | 0.709 | 0.703 |
| GS Num | 906k | 591k | 925k | 807k | |
| FPS | 1.16 | 1.3 | 1.15 | 1.20 | |
[…]
| SSIM | 0.754 | 0.780 | 0.590 | 0.708 | |
| SGS-SLAM | LPIPS | 0.289 | 0.318 | 0.463 | 0.357 |
| GS Num | 1201k | 1074k | 1049k | 1108k | |
| FPS | 0.25 | 0.25 | 0.24 | 0.25 | |
[…]
| SSIM | 0.874 | 0.861 | 0.823 | 0.853 | |
| GS4 (Ours) | LPIPS | 0.243 | 0.265 | 0.368 | 0.292 |
| GS Num | 182k | 167k | 185k | 178k | |
| FPS | 5.16 | 10.18 | 5.68 | 7.01 |
Appendix C Zero-Shot Results on TUM RGB-D
In Table C, we present the quantitative zero-shot results across three scenes from the TUM RGB-D dataset. TUM RGB-D provides ground truth camera trajectories, so we also report tracking performance. Our method achieves rendering performance comparable to that of all other GS-based SLAM approaches, while using significantly fewer Gaussians and offering much faster processing speed.
| Methods | Metrics | fr1desk | fr2xyz | fr3office | Avg |
| PSNR | 22.07 | 24.66 | 21.54 | 22.76 | |
| SSIM | 0.857 | 0.947 | 0.870 | 0.891 | |
| SplaTAM | LPIPS | 0.238 | 0.099 | 0.210 | 0.182 |
| ATE RMSE | 3.33 | 1.55 | 5.28 | 3.39 | |
| GS Num | 969k | 635k | 806k | 803k | |
| FPS | 0.32 | 0.05 | 0.25 | 0.20 | |
[…]
| SSIM | 0.715 | 0.795 | 0.797 | 0.769 | |
| RTG-SLAM | LPIPS | 0.438 | 0.353 | 0.394 | 0.395 |
| ATE RMSE | 1.66 | 0.38 | 1.13 | 1.06 | |
| GS Num | 236k | 84k | 273k | 198k | |
| FPS | 2.07 | 2.12 | 1.97 | 2.05 | |
[…]
| SSIM | 0.886 | 0.946 | 0.796 | 0.876 | |
| SGS-SLAM | LPIPS | 0.176 | 0.097 | 0.280 | 0.184 |
| ATE RMSE | 3.57 | 1.29 | 9.08 | 4.65 | |
| GS Num | 808k | 695k | 701k | 735k | |
| FPS | 0.25 | 0.10 | 0.20 | 0.18 | |
[…]
| SSIM | 0.855 | 0.908 | 0.840 | 0.868 | |
| GS4 (Ours) | LPIPS | 0.248 | 0.146 | 0.270 | 0.221 |
| ATE RMSE | 1.86 | 0.63 | 1.95 | 1.48 | |
| GS Num | 158k | 90k | 149k | 132k | |
| FPS | 1.34 | 7.26 | 2.78 | 3.79 |
Appendix D Zero-Shot Results on Replica
We present zero-shot visualizations on the Replica dataset in Fig. 6. As shown, our method produces more consistent renderings across various views. Note that our model used here is trained only on ScanNet.
Room1 |
SGS SLAM |
GS4 (Ours) |
Office3 |
SGS SLAM |
GS4 (Ours) |
Office4 |
SGS SLAM |
GS4 (Ours) |
| Ground-truth | RGB rendering | Semantic rendering |